OpenStudy (anonymous):

Determine a polynomial equation whose roots include the given values: 5,-5,1,-1

OpenStudy (anonymous):

\[(x ^{2}-5^{2})(x ^{2}-1)=0\]

OpenStudy (anonymous):

can you explain how you got there?

OpenStudy (anonymous):

now when you hav the roots to an equation as a. b etc... to get the equation you have to multiply by x-(root1)*x-(root2) and so on

OpenStudy (anonymous):

ahhhhh, ok cool!

OpenStudy (anonymous):

(x-a)*(x-b)

OpenStudy (anonymous):

for your values u get (x-5)(x+5)(x-1)(x+1)=0

OpenStudy (anonymous):

(a-b)*(a+b)=a^2-b^2

OpenStudy (anonymous):

expand and voila you get your equation

OpenStudy (anonymous):

take the roots as \[\alpha , \beta , \gamma , \delta\]

OpenStudy (anonymous):

then apply formula of quadratic eqn as u do for x^2 + (-b/a)x + c/a

OpenStudy (anonymous):

hey divya if u ever come back again physics_him@hotmail.com